# LitLLM: A Toolkit for Scientific Literature Review

### v1

@@ Abstract
Conducting literature reviews for scientific papers is essential for understanding research, its limitations, and building on existing work. It is a tedious task which makes an automatic literature review generator appealing. Unfortunately, many existing works that generate such reviews using Large Language Models (LLMs) have significant limitations. They tend to hallucinate—generate non-factual information—and ignore the latest research they have not been trained on. To address these limitations, we propose a toolkit that operates on Retrieval Augmented Generation (RAG) principles, specialized prompting and instructing techniques with the help of LLMs. Our system first initiates a web search to retrieve relevant papers by summarizing user-provided abstracts into keywords using an off-the-shelf LLM. Authors can enhance the search by supplementing it with relevant papers or keywords, contributing to a tailored retrieval process. Second, the system re-ranks the retrieved papers based on the user-provided abstract. Finally, the related work section is generated based on the re-ranked results and the abstract. There is a substantial reduction in time and effort for literature review compared to traditional methods, establishing our toolkit as an efficient alternative. Our open-source toolkit is accessible at https://github.com/shubhamagarwal92/LitLLM and Huggingface space (https://huggingface.co/spaces/shubhamagarwal92/LitLLM) with the video demo at https://youtu.be/E2ggOZBAFw0
@@
Shubham Agarwal1,2,3, Issam H. Laradji1,4, Laurent Charlin2,3,5, Christopher Pal1,2,5 1ServiceNow Research, 2Mila - Quebec AI Institute, 3HEC Montreal, Canada 4UBC, Vancouver, Canada, 5Canada CIFAR AI Chair Correspondence: shubham.agarwal@mila.quebec
@@ 1 Introduction
Figure 1: LitLLM on Huggingface Space. Our system works on the Retrieval Augmented Generation (RAG) principle to generate the literature review grounded in retrieved relevant papers. User needs to provide the abstract in the textbox (in purple) and press send to get the generated related work (in red). First, the abstract is summarized into keywords (Section 3.1), which are used to query a search engine. Retrieved results are re-ranked (in blue) using the Paper Re-Ranking module (Section 3.2), which is then used as context to generate the related work (Section 3.3). Users could also provide a sentence plan (in green) according to their preference to generate a concise, readily usable literature review (See Section 3.3.2).
@@ 3.1 Paper Retrieval Module
In our toolkit, we retrieve relevant papers using the Semantic Scholar API. Other platforms could be used, but the S2 Platform is well-adapted to this use case. It is a large-scale academic corpus comprising 200M+ metadata records across multiple research areas, providing information about papers’ metadata, authors, paper embedding, etc. The Recommendations API also provides relevant papers similar to any seed paper. Figure 3 shows our system’s different strategies. We describe these three settings that we use to search for references:
@@ 3.2 Paper Re-Ranking Module
Recent efforts have explored the application of proprietary LLMs for ranking  Sun et al. (2023); Ma et al. (2023) as well as open-source models like  Pradeep et al. (2023a, b). These approaches provide a combined list of passages directly as input to the model and retrieve the re-ordered ranking list  Zhang et al. (2023). Typically, a retriever first filters top-k potential candidates, which are then re-ranked by an LLM to provide the final output list. In our work, we use the instructional permutation generation approach Sun et al. (2023) where the model is prompted to generate a permutation of the different papers in descending order based on the relevance to the user-provided abstract, thus producing an ordered list of preferences against providing intermediate scores. Figure 5 showcases the prompt we used for LLM-based re-ranking.
@@ 4 Implementation Details
We build our system using Gradio Abid et al. (2019), which provides a nice interface to quickly and efficiently build system demos. Our user interface is also available at HuggingFace Space [^5] https://huggingface.co/spaces/shubhamagarwal92/LitLLM. We query the Semantic Scholar API available through the Semantic Scholar Open Data Platform (Lo et al., 2020; Kinney et al., 2023) to search for the relevant papers. Specifically, we use the Academic Graph [^6] https://api.semanticscholar.org/api-docs/graph and Recommendations [^7] https://api.semanticscholar.org/api-docs/recommendations API endpoint. In this work, we use OpenAI API [^8] https://platform.openai.com/docs/guides/gpt to generate results for LLM using GPT-3.5-turbo and GPT-4 model. At the same time, our modular pipeline allows using any LLM (proprietary or open-sourced) for different components. We also allow the end-user to sort the retrieved papers by relevance (default S2 results), citation count, or year.
@@ 6 Conclusion and Future Work
The field of multimodal research has seen significant advancements in recent years, with the development of models such as Flamingo and DALL-E that focus on image captioning and text-to-image generation respectively. However, the question of whether these models can understand each other and work in harmony is a topic of interest. In this context, the work of [1] presents the Contrastive Captioner (CoCa), a model that combines contrastive loss and captioning loss to pretrain an image-text encoder-decoder foundation model. This model, while efficient, does not fully address the interaction between image and text modalities at a fine-grained level.
The work of [2] addresses this gap by proposing a jointly masked multimodal modeling method that focuses on fine-grained multimodal representations. This method performs joint masking on image-text input and integrates both implicit and explicit targets for the masked signals to recover. However, this approach, while effective, does not fully address the challenges of intra-modal reasoning and cross-modal alignment that are inherent in image-text retrieval.
The work of [3] presents a novel modality interaction modeling network based on the routing mechanism, which is the first unified and dynamic multimodal interaction framework towards image-text retrieval. This model can dynamically learn different activated paths for different data, providing a more flexible approach to modality interaction. However, the design of interaction patterns in this model still relies heavily on expert experience and empirical feedback, which may limit its applicability in different contexts. The work of [4] introduces the Wikipedia-based Image Text (WIT) Dataset, a large-scale dataset for multimodal, multilingual learning. This dataset, while providing a rich resource for multimodal learning, does not directly address the question of how different models can understand each other and work together.
In light of these previous works, our study proposes a reconstruction task where Flamingo generates a description for a given image and DALL-E uses this description as input to synthesize a new image. We argue that these models understand each other if the generated image is similar to the given image. This approach allows us to study the relationship between the quality of the image reconstruction and that of the text generation, and to propose a unified framework to finetune the text-to-image and image-to-text models. Our extensive experiments validate our findings and demonstrate the effectiveness of our proposed unified framework.

### v2

@@ Abstract
Conducting literature reviews for scientific papers is essential for understanding research, its limitations, and building on existing work. It is a tedious task which makes an automatic literature review generator appealing. Unfortunately, many existing works that generate such reviews using Large Language Models (LLMs) have significant limitations. They tend to hallucinate—generate non-factual information—and ignore the latest research they have not been trained on. To address these limitations, we propose a toolkit that operates on Retrieval Augmented Generation (RAG) principles, specialized prompting and instructing techniques with the help of LLMs. Our system first initiates a web search to retrieve relevant papers by summarizing user-provided abstracts into keywords using an off-the-shelf LLM. Authors can enhance the search by supplementing it with relevant papers or keywords, contributing to a tailored retrieval process. Second, the system re-ranks the retrieved papers based on the user-provided abstract. Finally, the related work section is generated based on the re-ranked results and the abstract. There is a substantial reduction in time and effort for literature review compared to traditional methods, establishing our toolkit as an efficient alternative. Our project page including the demo and toolkit can be accessed here: https://litllm.github.io.
@@
Shubham Agarwal*1,2,3, Gaurav Sahu*1, Abhay Puri*1, Issam H. Laradji1,4, Krishnamurthy DJ Dvijotham1, Jason Stanley1, Laurent Charlin2,3,5, Christopher Pal1,2,5 1ServiceNow Research, 2Mila - Quebec AI Institute, 3HEC Montreal, Canada 4UBC, Vancouver, Canada, 5Canada CIFAR AI Chair, 6University of Waterloo Correspondence: {shubham.agarwal, gaurav.sahu}@mila.quebec Journal version: https://openreview.net/forum?id=heeJqQXKg7 thanks: Equal contribution.
@@ 1 Introduction
Figure 1: LitLLM Interface. Our system works on the Retrieval Augmented Generation (RAG) principle to generate the literature review grounded in retrieved relevant papers. User needs to provide the abstract in the textbox (in purple) and press send to get the generated related work (in red). First, the abstract is summarized into keywords (Section 3.1), which are used to query a search engine. Retrieved results are re-ranked (in blue) using the Paper Re-Ranking module (Section 3.2), which is then used as context to generate the related work (Section 3.3). Users could also provide a sentence plan (in green) according to their preference to generate a concise, readily usable literature review (See Section 3.3.2).
@@ 3.1 Paper Retrieval Module
In our toolkit, we retrieve relevant papers using the Semantic Scholar API Kinney et al. (2023) and OpenAlex API Priem et al. (2022). Other platforms could be used, but the S2 and OpenAlex platforms are well-adapted to this use case. Combined, they constitue a large-scale academic corpus comprising 300M+ metadata records across multiple research areas, providing information about papers’ metadata, authors, paper embedding, etc. The S2 Recommendations API also provides relevant papers similar to any seed paper. Figure 3 shows our system’s different strategies. We describe these three settings that we use to search for references:
@@ 3.2 Paper Re-Ranking Module
Recent efforts have explored the application of proprietary LLMs for ranking  Sun et al. (2023); Ma et al. (2023) as well as open-source models like  Pradeep et al. (2023a, b). These approaches provide a combined list of passages directly as input to the model and retrieve the re-ordered ranking list  Zhang et al. (2023). Typically, a retriever first filters top-k potential candidates, which are then re-ranked by an LLM to provide the final output list. In our work, we explore the instructional permutation generation approach Sun et al. (2023) where the model is prompted to generate a permutation of the different papers in descending order based on the relevance to the user-provided abstract, and debate-ranking with attribution Rahaman et al. (2024); Agarwal et al. (2024) where an LLM is prompted to (1) generate arguments for and against including the candidate paper and (2) output a final probability of including the candidate based on the arguments. Figure 5 showcases the prompt we used for LLM-based re-ranking.
@@ 4 Implementation Details
We build our system using React, which provides a nice interface to quickly and efficiently build system demos. We query the Semantic Scholar API available through the Semantic Scholar Open Data Platform (Lo et al., 2020; Kinney et al., 2023) to search for the relevant papers. Specifically, we use the Academic Graph [^5] https://api.semanticscholar.org/api-docs/graph and Recommendations [^6] https://api.semanticscholar.org/api-docs/recommendations API endpoint. For OpenAlex, we query the search endpoint.  [^7] https://api.openalex.org/works In this work, we use OpenAI API [^8] https://platform.openai.com/docs/guides/gpt to generate results for LLM using GPT-3.5-turbo and GPT-4 model. At the same time, our modular pipeline allows using any LLM (proprietary or open-sourced) for different components. We also allow the end-user to sort the retrieved papers by relevance (default S2 results), citation count, or year.
@@ 6 Conclusion and Future Work
The field of multimodal research has seen significant advancements in recent years, with the development of models such as Flamingo and DALL-E that focus on image captioning and text-to-image generation respectively. However, the question of whether these models can understand each other and work in harmony is a topic of interest. In this context, the work of [1] presents the Contrastive Captioner (CoCa), a model that combines contrastive loss and captioning loss to pretrain an image-text encoder-decoder foundation model. This model, while efficient, does not fully address the interaction between image and text modalities at a fine-grained level. The work of [2] addresses this gap by proposing a jointly masked multimodal modeling method that focuses on fine-grained multimodal representations. This method performs joint masking on image-text input and integrates both implicit and explicit targets for the masked signals to recover. However, this approach, while effective, does not fully address the challenges of intra-modal reasoning and cross-modal alignment that are inherent in image-text retrieval. The work of [3] presents a novel modality interaction modeling network based on the routing mechanism, which is the first unified and dynamic multimodal interaction framework towards image-text retrieval. This model can dynamically learn different activated paths for different data, providing a more flexible approach to modality interaction. However, the design of interaction patterns in this model still relies heavily on expert experience and empirical feedback, which may limit its applicability in different contexts. The work of [4] introduces the Wikipedia-based Image Text (WIT) Dataset, a large-scale dataset for multimodal, multilingual learning. This dataset, while providing a rich resource for multimodal learning, does not directly address the question of how different models can understand each other and work together. In light of these previous works, our study proposes a reconstruction task where Flamingo generates a description for a given image and DALL-E uses this description as input to synthesize a new image. We argue that these models understand each other if the generated image is similar to the given image. This approach allows us to study the relationship between the quality of the image reconstruction and that of the text generation, and to propose a unified framework to finetune the text-to-image and image-to-text models. Our extensive experiments validate our findings and demonstrate the effectiveness of our proposed unified framework.